For each question determine if the situation refers to a mean or a proportion and indicate it by highlighting your choice then, solve each problem accordingly.
Each question is worth 10 points, 2 points for the correct choice and 8 points for the correct answer.
1. Find the sample size needed to estimate the percentage of republicans among registered voters in California. Use a 0.03 margin of error; use a confidence level of 90%.
Mean Proportion

Answer by stanbon(75887)

Find the sample size needed to estimate the percentage of republicans among registered voters in California. Use a 0.03 margin of error; use a confidence level of 90%.
----
Percentage is a proportion.
-------
Since E = z*sqrt(pq/n)
--------
n = (z/E)^2*p*q = (1.645/0.03)^2*(1/2)(1/2) = 752 when rounded up
-----------
Cheers,
Stan H.
-----------